OpenStudy (anonymous):

The equation of a circle whose center is at (4, 0) and radius is length 2√(3).

OpenStudy (dtan5457):

\[(x-h)^2+(y-k)^2=r^2\] h,k is your center point. You can square your radius to make it a whole number

OpenStudy (anonymous):

so would it be (x-2)^2+(y-5)^2=?

OpenStudy (anonymous):

i got 45

OpenStudy (dtan5457):

You don't really solve for the equation..look